OpenStudy (anonymous):

rewrite as a split function: f(x) = |3x - x^2|

OpenStudy (anonymous):

when, 3x-x^2 > 0, f(x) = 3x-x^2 when 3x-x^2 < 0 , f(x) = -(3x-x^2) = x^2 - 3x

OpenStudy (anonymous):

simplify the conditions 3x-x^2 > 0 and 3x-x^2 < 0 further to get the ranges of x for both cases.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its been a pleasure :)